Reports 1-1 of 1 Clear search Modify search
AdV-COM (AdV commissioning (1st part) )
bersanetti - 13:04 Saturday 23 September 2023 (61759) Print this report
Lock (acquisition) stability

I gave a look at the unlocks from this morning, both during lock acquisition and in the "steady" states; besides a couple of them, they fall in three categories:

  • unlocks during the CARM handoff to the IMC: long standing story, they have been quite reduced in frequency after the last change in CARM offset at the time of the handoff, but it looks like that bad weather (as in wind and sea activity, which today are both annoying) has an impact on those;
  • unlocks at apparently steady state in CARM_NULL_1F: "apparently" as they happen just a few seconds after getting there; I could find one definite instance, and two other suspect ones, of the fact that they are induced by the engagement of the CARM2RFC loop, which either saturates (Figure 1) and/or pushes a lot with "fast" transients +-5/6 V (Figures 2 and 3); for the time being I tried using a different CARM filter (1 success out of 2 at the time of writing), I will observe during the day and maybe change strategy (lower gain to act more as a drift control to avoid fast/big corrections);
  • unlocks in LOW_NOISE_2: a couple of them happened (Figures 4 to 7, two unlocks) with a quickly diverging 55 Hz oscillation, most probably induced by the diagonalization process which abruptly changes sign to a couple of contributions; I disabled the re-engagement of PyDiag in LOW_NOISE_2 and will observe behaviour/performance.
Images attached to this report
Comments to this report:
bersanetti - 19:51 Saturday 23 September 2023 (61760) Print this report

The situation seems improved up to CARM Null (but the wind also went down); instead, after DC Readout the B1 PDs keep saturating (Figures 1 and 2), also after the LOW_NOISE_1 transition (where they are forced to use DC), so during the LOW_NOISE_2 transition and in steady state (as in the 1 minute lock). As a test I completely disabled the saturation switch, so it remains forced to 1 and the B1 PDs will keep using the DC and not the Audio information. First test shortly in the current acquisition.

FOLLOW-UP: also this did not work, B1 consistently on DC but glitches on DARM and correction saturation. SR alignment gets spoiled (because of the DARM glitches?), and I don't know how the OMC reacts to glitches on B1_PD2 (Figure 3); DARM_UGF not informative as coherence too low, DARM_OG really low.

Images attached to this comment
bersanetti, gouaty, magazzu - 22:06 Saturday 23 September 2023 (61763) Print this report

The alternative CARM filter, while helping in bad weather keeping the lock at CARM_NULL_1F, is too rough and corrects too much for LOW_NOISE operation, and it caused at least part of the problems there. Switching back in-lock to the old filter was beneficial to LOW_NOISE_2 )no saturation) and, after a while (low Range segment) we put back the automatic switch to the Audio channels, which brought back the ITF close to nominal. Given that the weather is also better, I restored both the filter and the automatic switch in the automation. If in bad weather we will still face problems a two-stages CARM strategy will be implemented (after a revision of the alternative filter).

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×